# Predictive Analysis for Optimizing Port Operations

### v1

@@ 4 Results
Figure 8: Feature importance for of Delay time classification at port
Figure 9: Confusion matrix on test data for Delay time classification at port
@@ 5 Conclusion
In conclusion, the development of a comprehensive prediction and classification solution for estimating vessel stay and Delay times in maritime logistics represents a significant step forward in addressing the complexities of port operations. By providing tools to accurately predict and classify vessel behavior, this research contributes to more efficient and effective decision-making in port management. The wide-ranging applications of the proposed solution, from ship scheduling to service delay predictions to risk assessment, demonstrate its potential to enhance overall port performance and optimize supply chain operations. Additionally, the incorporation of feature analysis enriches our understanding of the factors influencing maritime logistics, where predictions are influenced by Berth and Cargo Characteristics features and classification tasks are impacted by Geographical features (Total time) and operation characteristic features (Delay time). This offers insights that can be leveraged to further improve the efficiency of port operations and make them more resilient. Since Berth plays a key role in prediction of Total time and Delay time, using a Top-Bottom approach for developing individual model for ports and then berths would be beneficial in understanding how these ports and berths can be made more efficient.
@@
Focus on refining the prediction and classification models to account for additional variables and complexities in maritime logistics is a key area of future work. This could involve integrating real-time data streams, such as weather forecasts, vessel tracking, other vessel information, to enhance the accuracy and responsiveness of the models. Furthermore, exploring the application of more advanced techniques, such as deep learning and reinforcement learning, may offer new avenues for improving the predictive capabilities of the models. Additionally, efforts to expand the scope of the research beyond individual ports to encompass broader regional or global maritime networks could provide valuable insights into optimizing port operations within larger logistical contexts. Overall, continued research in this area has the potential to drive significant advancements in maritime logistics, leading to more efficient and resilient port operations in the future.

### v2

@@ 4 Results
In the classification task, the performance of various Machine Learning methods for (Multi-Class) classification of Total Time is depicted in Table 3. A consistent trend is observed where tree-based methods outperform statistical approaches as well as deep learning methods again, potentially due to the complex nature of the data and high number of categorical features. Specifically, Extreme Gradient Boosting (XGBoost) demonstrates the best performance, achieving the highest Accuracy (0.8424) and superior values across other metrics compared to alternative methods. This is 10% higher than the result in [Abreu et al., 2022]. The test set results further confirm XGBoost’s efficacy with an Accuracy of 0.8415 and other metrics exhibiting similarly favorable performance, derived from the confusion matrix in Fig. 6(a). Also, we observe that is it easier to detect the outer classes, Class 0 (Low) and Class 3 (Very High), compared to the other two.
Additionally, Fig. 6(b) showcases the key features influencing Total Time classification. Notably, the Geographical Region (North), Port Complex (Manaus), Type of Mooring Navigation (inside) emerge as the most influential features, followed by factors such as City, Mooring port, Inland Transportation route, Type of Mooring Navigation (Long Haul), Berth, etc. The significance of the geographical features is related to the huge coastline of Brazil and the complex dynamics of port operations. The SHAP analysis is illustrated in Fig. 7. Specifically, Fig. 7(a) demonstrates that features such as Berth, Destination, Moved Value, and Month have the most significant impact on the outcomes of the XGBoost model. Each sample will have 4 Decision and Force plots, one for each class, since the sample considered has a class of 1, we shared the plots of Class 1. Fig. 7(b) and Fig. 7(c) present the Decision and Force plots for Class 1 for a particular sample with a true Total Time class as 1 (Medium). The model output is 1.64, which is the highest among the four classes and hence this sample is correctly classified as class 1. These plots reveal that features like Berth, Destination, and Moved Value play a crucial role in producing the accurate classification for this specific sample. Reducing the Moved Value for this sample should push it towards Class 0 (Low), which intuitively makes sense but may not be feasible.
(a) Confusion matrix on test data for Total Time classification
(b) Random Forest Feature importance for of Total Time classification
Figure 6: Random Forest inference plots for Total Time classification
(a) SHAP feature analysis plot for Total Time classification
(b) SHAP Decision plot for Total Time classification of a sample observation
(c) SHAP force plot of a sample observation for Class 1
Figure 7: SHAP Analysis for Total Time classification
Table 4 presents the binary classification performance on Delay Time. The tree-based methods are again outperforming other approaches. XGBoost gives the best performance by achieving the highest Accuracy (0.9511) and best values across the different metrics compared to other methods. XGBoost achieves an Accuracy of 0.9505 on the test set and other metrics showing close performance, as illustrated by the confusion matrix in Fig. 8(a).
(a) Confusion matrix on test data for Delay Time classification
(b) Random Forest Feature importance for of Delay Time classification
Figure 8: Random Forest inference plots for Delay Time classification
(a) SHAP feature analysis plot for Delay Class
(b) SHAP Decision plot for Delay Class of a sample observation
(c) SHAP force plot for Delay Class of a sample observation
Figure 9: SHAP Analysis for Delay Class
Furthermore, Fig. 8(b) highlights the key features influencing Delay Time classification, factors such as Flag Long Course, Flag Cabotage, Owner’s Nationality, and Type of Mooring Navigation (Long Haul) emerge as the most influential. This is followed by variables like Port Complex, Load Operation Type, ST Cargo Nature, and Month. Features related to operation characteristics are having the most impact on the Delay class. Fig. 9 shows the SHAP analysis for Delay Class, where Fig. 9(a) indicates that the most significant impact on the XGBoost output is by the features Berth, Source, destination, and Nationality. Figs. 9(b) and 9(c) show the Decision plot and Force plot for a specific sample, where the actual delay class is 0. The SHAP output is -5.08, indicating the class is class 0 since it has a low value, with features such as Month, Berth, and Flag Cabotage having the most influence on this sample. Since this sample is Class 0, meaning no delay, we don’t want to change any feature in this case.
@@ 5 Conclusion
In conclusion, this development of a comprehensive prediction and classification solution for estimating vessel Stay (Total) Time and Delay Time in maritime logistics represents a significant step forward in addressing the complexities of port operations. Our study highlights the use and importance of predictive analysis for port operations using Brazilian port data. This can easily be extended to benefit port operations worldwide, offering opportunities for improved efficiency and effectiveness in maritime logistics on a global scale. Notably, tree-based methods exhibit the best performance, particularly in predicting Total Time and Delay Class. By providing tools to accurately predict and classify vessel behavior, this research contributes greatly to improving our understanding of various methods, which can be used to make more efficient and effective predictive analytics based decision-making in port management. The proposed solution has wide-ranging applications and integration into different aspects of maritime logistics, from ship scheduling to service delay predictions to risk assessment, demonstrating its potential to enhance overall port performance and optimize supply chain operations.
Additionally, the incorporation of feature analysis through feature importance enriches our understanding of the factors influencing maritime logistics, where predictions are influenced by Berth and Cargo Characteristics features while classification tasks are impacted by Geographical features (Total Time) and operation characteristic features (Delay Time). This offers insights that can be leveraged to further improve the efficiency of port operations and make them more resilient. Since Berth plays a key role in the prediction of Total Time and Delay Time, using a Top-Bottom approach for developing individual models for ports and then berths would be beneficial in understanding how these ports and berths can be made more efficient. SHAP (SHapley Additive exPlanations) analysis plays a crucial role in this process by helping us understand how these features contribute to the outcomes, providing deeper insights into each individual case. This allows for the identification of specific mitigation steps that can be taken accordingly. We observe that Berth, Source, and Destination, and Month are usually coming out as key features impacting outcomes. Learning from this, we can suggest an optimal Berth, Source, Destination combination and an ideal travel month if operationally feasible. SHAP helps us address the predictive analytics of port operations at both the global and local levels, ensuring that the models are both comprehensive and precise. Overall, this comprehensive approach, enhanced by Feature Importance and SHAP analysis, contributes to a more nuanced understanding of port logistics and supports the development of more effective and resilient port management strategies.
@@
Deep learning methods do not currently perform well due to the abundance of categorical data. In the future, we plan to address this issue by exploring alternative methods for handling categorical data, such as binary encoding, deep embedding (Cat2Vec), and other techniques, to see if they can improve the performance of deep learning models. We would like to focus on refining the prediction and classification models to account for additional variables and complexities in maritime logistics as a key area of future work. It could involve integrating real-time data streams, such as weather forecasts, vessel tracking, and other vessel information, to enhance the accuracy and responsiveness of the models. These expanded features will lead to enhanced feature analysis as well. Furthermore, exploring the application of more advanced techniques may offer new avenues for improving the modeling capabilities. Additionally, efforts to expand the scope of the research beyond individual ports to encompass broader regional or global maritime networks could provide valuable insights into optimizing port operations within larger logistical contexts. Overall, continued research in this area has the potential to drive significant advancements in maritime logistics, leading to more efficient and resilient port operations in the future.